OpenStudy (anonymous):

Find the slope of a line passing through. (0, -5 and (-4 ,7)

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow {y_{2} - y_{1} \over x_{2} - x_{1}} }\) \(\Large \color{MidnightBlue}{\Rightarrow {7 + 5 \over -4} }\) \(\Large \color{MidnightBlue}{\Rightarrow {12 \over -4} }\) \(\Large \color{MidnightBlue}{\Rightarrow -3 }\)

OpenStudy (anonymous):

m = 7 -(-5) / -4 - 0 m = 12 / -4 m = -3/1 m = -3

OpenStudy (anonymous):

@MayoriShira and i think your a owlet...no offense.

OpenStudy (anonymous):

@MayoriShira i think your a f a g no offense

OpenStudy (campbell_st):

the gradient or slope is \[m = \frac{(y _{2}-y _{1})}{x _{2} -x _{1}}\] \[m = \frac{7 + 5}{-4 - 0} = \frac{12}{-4} =-3\]

OpenStudy (anonymous):

uh dude chill =A= sorry

Parth (parthkohli):